In a proton accelerator used in elementary particle physics experiments, the trajectories of protons are controlled by bending magnets that produce a magnetic field of 2.80 T.
What is the magnetic-field energy in a 11.0 cm^3 volume of space where = 2.80 T?
Answer #1
Concepts and reason

The concepts used to solve this problem are energy density in magnetic field and magnetic field energy.

Initially, use the relationship between magnetic field, energy density and permeability to find the energy density in magnetic field.

Finally, use the relationship between volumes of space, energy density in magnetic field to find the magnetic field energy.

Fundamentals

The amount of energy saved in a system per unit volume is explained as energy density.

The expression energy density in magnetic field is given below:

μg
2μ,

Here, magnetic field is , energy density in magnetic field is , and permeability is .

The expression for magnetic field energy of a given volume of space is given below:

E =(up)(v)

Here, magnetic field energy is , and volume is .

The expression energy density in magnetic field is given below:

μg
2μ,

Substitute 47x10-7Tm/A
for and 2.80T
for .

Us
(2.80T)
2(41 x10-?Tm/A)
=0.3119x10J/m

The expression for magnetic field energy of a given volume of space is given below:

E =(up)(v)

Substitute 0.3119x10J/m
for and 11.0cm
for .

E =(0.31192.10* Jm) (11.0cm (i om
=(0.3119x10J/m) (11.0x10 m)
= 34.309)

Ans:

Thus, the magnetic field energy of a given volume of space is 34.309J
.